Ellis issues 8.5%, five-year bonds dated January 1, 2017, with a $540,000 par value. The bonds pay interest on June 30 and December 31 and are issued at a price of $597,579. The annual market rate is 6% on the issue date.

1. Complete the below table to calculate the total bond interest expense over the bonds' life.
2. Prepare a straight-line amortization table for the bonds’ life.
3. Prepare the journal entries to record the first two interest payments.
Answer #1
1
Total interest expense over life of bonds
10 payments of $ 22950 229500
Par value at maturity 540000
Total repaid 769500
Less: Amount borrowed 597579
Total bond interest expense 171921
2
Semiannual Interest period end Unamortized Premium Carrying value
01/01/2017 57579 597579
06/30/2017 51821 591821
12/31/2017 46063 586063
06/30/2018 40305 580305
12/31/2018 34547 574547
06/30/2019 28789 568789
12/31/2019 23031 563031
06/30/2020 17273 557273
12/31/2020 11515 551515
06/30/2021 5757 545757
12/31/2021 0 540000
3
Debit Credit
June 30,2017 Interest expense 17192
Premium on Bonds payable 5758 =(597579-540000)/10
       Cash 22950 =540000*8.5%/2
Debit Credit
December 31,2017 Interest expense 17192
Premium on Bonds payable 5758 =(597579-540000)/10
       Cash 22950 =540000*8.5%/2